# Dynamical Phase Transitions Across Slow and Fast Regimes in a Two-Tone Driven Duffing Resonator

### v1

@@ VI Phase Transition Boundaries
Beyond the specific case studied here, other combinations of thresholds and trajectories are possible depending on the initial starting stationary state and the chosen primary tone detuning. For instance, starting from the higher branch, the system will leave the higher branch towards the lower branch if ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},-$}}\leq F_{hb}$ and will reach the lower branch if ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},-$}}\leq F_{hb}$. These additional scenarios are systematically summarized in Table 1, which maps the inequalities between ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},+$}}$, ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{h},-$}}$, $F_{lb}$, and $F_{hb}$ onto the corresponding dynamical outcomes. This overview highlights that the observed behaviour is just one instance within a broader set of possible instability regimes, underscoring the richness of the system’s driven dynamics, cf. Appendix B for additional examples. There, too, depending on the particular parameters used, the thresholds obtained from our model can deviate from the boundary observed in simulations. These deviations also trace back to the simplifying assumption of using a two-tone driven linear response to approximate the two-tone driven Duffing resonator’s nonlinear response.
VII Conclusion/Outlook
In this work, we established an analytical framework to describe dynamical phase transitions in a Duffing resonator under bichromatic driving. We reveal two regimes: a slow-beating one, where the secondary tone slowly modulates the main drive and can push the system past bifurcations, and a fast-modulation one. Our analysis shows that even a weak secondary tone can profoundly reshape the dynamics, inducing transitions between coexisting attractors that cannot be explained by perturbative treatments of the secondary tone. This provides a qualitative yet predictive tool to detect and categorize different types of dynamical phase transitions in two-tone driven nonlinear systems.
In contrast to earlier studies that surveyed specific parameters from the perspective of application in sensing [29] and emphasized cascades to chaos [30], we focused on rigorous analysis in a regime spanning a wide range of parameters. This allowed us to pinpoint the mechanism by which inter-attractor transitions occur and to connect the onset of switching to the resonance properties of the target state. Our linear-power-based framework leads to threshold conditions that describe inter-attractor trajectories across nanomechanical [32], optical [37], and superconducting resonator [54, 38] platforms.
Looking ahead, our framework provides a foundation for systematic extensions. It can incorporate additional tones, developing a multifrequency expansion beyond the heuristic approximations used here, and analytically clarifying the route to chaos. Moreover, we could account for stochastic or quantum-activation–driven transitions [55], and offer insight into quantum control and qubit operation in strongly nonlinear Kerr resonators. More broadly, our results establish multi-tone driving as a controlled route to engineer dynamical phase transitions, with direct impact on Kerr comb generation and nonlinear resonator control [56, 57, 39].
@@ Acknowledgements
We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG) through project numbers 449653034, 521530974, and 545605411, as well as via SFB 1432 (project number 425217212). We also acknowledge support from the Swiss National Science Foundation (SNSF) through the Sinergia Grant No. CRSII5_206008/1. J.d.P. acknowledges support from the Ramón y Cajal program (RYC2023-043827-I).
@@ Appendix B Exploration of more dynamical cases
Figure 7: The average amplitude phase diagram with threshold lines [cf. Fig. 5(c)], for an initial point starting from the higher branch, for $F_{1}\sim F_{lb}$. The parameters used are $\Omega_{0}=2\pi$$\Delta_{10}=-9.02\times 10^{-4}$, $\Gamma=6.2\times 10^{-4}$, and $F_{1}=6.01\times 10^{-4}$.
@@ Appendix C Experimental platform
$\ddot{x}+\Omega_{n}^{2}x+\Gamma\dot{x}=F_{\mathrm{dr}}(t),$ (26)
where $\Omega_{n}$ and $\Gamma$ are the resonance frequency and the damping rate of the $n$-th mode, respectively, and $F_{\mathrm{dr}}(t)=KU(t)$ is the applied drive in units of $\mathrm{V}$\,$\mathrm{s}$^{-2}$. $K$ is a conversion factor expressed in units of $\mathrm{s}$^{-2}$. When the strength of the external driving force increases, the resonator behavior starts to deviate from the standard damped harmonic oscillator. As we enter such large amplitude oscillation limit, to describe the system response to an external driving force we need to use the Duffing equation:
$\ddot{x}+\Omega_{n}^{2}x+\Gamma\dot{x}+\alpha x^{3}=F_{\mathrm{dr}}e^{i(\Omega_{\mathrm{dr}}t+\theta)},$ (27)
@@
In this work, we only focus on the lowest mechanical mode of the resonator. The mechanical parameters for this particular device have been characterized in a previous work [31] by measuring the amplitude ($A$) and phase ($\phi$) response to an external drive in both the weak and strong drive limit. Using a lock-in amplifier, we apply an external driving force $\Omega_{dr}$ to the system, varying its value across the resonance $\Omega_{0}$. With the same instrument, we simultaneously measure $A$ and $\phi$ in the frame rotating at $\Omega_{dr}$. The measured mechanical parameters are $\Omega_{0}/2\pi=1.11\,$\mathrm{M}\mathrm{H}\mathrm{z}$\$, $\Gamma/2\pi=108\,$\mathrm{H}\mathrm{z}$\$, $\alpha=-1.89\,$\mathrm{V}$^{-2}\,\mathrm{s}^{-2}$ and $K\approx 1\times 10^{7}s^{-2}$.
To generate the phase diagram in Fig. 5(a), we measure the resonator amplitude $A$ as a function of time when driven with two different tones $F_{1}e^{i(\Omega_{1}t)}$ and $F_{2}e^{i(\Omega_{2}t))}$. We use the lock-in amplifier to apply the two driving tones and readout the amplitude in the frame rotating at $\Omega_{1}$. After fixing $F_{1}/K=140\mathrm{mV}$ and $\Delta_{10}/2\pi\approx-158\,{Hz}$, we systematically probe the response of the resonator while varying $\Delta_{21}=\Omega_{2}-\Omega_{1}$ and $h$. For each combination, we extract the average value $\bar{A}$ for an interval of $5\,$\mathrm{s}$\$ making sure that for each combination the system is initialized in the lower stable solution of the Duffing curve before turning on the second drive. An example of measured $A(t)$ in the two limits of $\Delta_{21}<\Gamma$ and $\Delta_{21}\approx\Gamma$ are shown in Fig. 10.

### v2

@@ VI Phase Transition Boundaries
Beyond the specific case studied here, other combinations of thresholds and trajectories are possible depending on the initial starting stationary state and the chosen primary tone detuning. For instance, starting from the higher branch, the system will leave the higher branch towards the lower branch if ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}\leq F_{\mathrm{hb}}$ and will reach the lower branch if ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{0},-$}}\leq F_{\mathrm{hb}}$. These additional scenarios are systematically summarized in Table 1, which maps the inequalities between ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},+$}}$, ${F_{\mathrm{eff}}}_{\raisebox{-0.90417pt}{$\scriptscriptstyle\Omega_{\mathrm{h}},-$}}$, $F_{\mathrm{lb}}$, and $F_{\mathrm{hb}}$ onto the corresponding dynamical outcomes. This overview highlights that the observed behaviour is just one instance within a broader set of possible instability regimes, underscoring the richness of the system’s driven dynamics, cf. Appendix B for additional examples. There, too, depending on the particular parameters used, the transition thresholds obtained from our model can deviate from the boundary observed in simulations. These deviations also trace back to the simplifying assumption of using a two-tone driven linear response to approximate the two-tone driven Duffing resonator’s nonlinear response.
VII Extension of the Ansatz
Figure 6: Fourier spectra of the trajectories from Fig. 3. The peaks at $\omega=0$ (blue) and $\omega=\Delta_{21}$ (pink) correspond to the response components at the primary tone $\Omega_{1}$ and secondary tone $\Omega_{2}$, respectively, in the $\Omega_{1}$-rotating frame. Grey markers indicate the stationary state amplitudes $A_{1}$ and $A_{2}$ predicted by the multifrequency HBM ansatz [Eqs. (22)]. The other parameters used are $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, $F_{1}=4.81\times 10^{-4}$, and $\Delta_{10}=-9.02\times 10^{-4}$.
In Sections V and VI, we analyzed how the secondary tone modifies the stationary response of the primary tone. The analysis stemmed from a quasistatic picture which focused on the secondary tone’s influence on the response of the primary tone and ignored nonlinearity-induced interaction between the responses at the two tones. However, the complex beating patterns observed in Fig. 3 suggest that the interplay between the two drives generates dynamics that the single-tone ansatz [Eq. (3)] cannot capture. To clarify the spectral content of these dynamics, we compute the Fourier transform of the trajectories in Fig. 3. Tracking different Fourier components allows the definition of a "stationary state", even though the dynamics appear non-stationary in a single-frequency rotating frame. As shown in Fig. 6, the spectrum reveals a significant component at the detuning frequency $\omega=\Delta_{21}$, corresponding to the $\Omega_{2}$ response in the rotating frame. To capture this, we extend the single-tone ansatz [Eq. (3)] to a superposition of two rotating frames, treating both driving tones on an equal footing:
$x(t)=\sum_{i=1,2}X_{i}(T)\cos{(\Omega_{i}t)}-Y_{i}(T)\sin{(\Omega_{i}t)}\,$, (21)
where $X_{i}$, $Y_{i}$ are the quadratures corresponding to the response at $\Omega_{i}$.
Figure 7: Stationary states obtained from HBM using a multifrequency ansatz [cf. Eqs. (22)] for fixed (a) $\Delta_{21}=0.85\Gamma$ and (b) $\Delta_{21}=-0.85\Gamma$ as a function of $h$. The coloured lines in the top (bottom) panel represent $A_{1}$ at $\Omega_{1}$ ($A_{2}$ at $\Omega_{2}$) and the grey lines represent $A_{2}$ ($A_{1}$). The colours identify the different branches: with respect to $A_{1}$, blue (pink) refers to the low- (high-) amplitude response at the main drive tone; the corresponding amplitude $A_{2}$ is colored accordingly. The solid lines are stable branches and the dashed lines are saddles. The dotted vertical lines correspond to the (black, blue, and pink) transition boundaries shown in Fig. 5 for the respective $\Delta_{21}$. Other parameters are $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, $F_{1}=4.81\times 10^{-4}$, and $\Delta_{10}=-9.02\times 10^{-4}$.
We use the multifrequency ansatz (21) and apply the HBM, similar to Section IV. In the far-detuned regime ($\Delta_{21}\gg\Gamma$), the system’s response time $\sim 2\pi/\Gamma$ is much longer than the beating period $2\pi/|\Delta_{21}|$. Thus, we can assume that the quadratures evolve on a timescale $T$ much slower than the driving periods, $T_{i}=2\pi/\Omega_{i}$. This yields four coupled first-order differential equations in $X_{i}$, $Y_{i}$:
$\dot{X}_{1}$ $=-\frac{\Gamma}{2}X_{1}+\left[\frac{\tilde{\Delta}^{2}_{01}}{2\Omega_{1}}+\frac{3\alpha}{4\Omega_{1}}(\frac{A_{1}^{2}}{2}+A_{2}^{2})\right]Y_{1}+\frac{F_{1}\sin{\theta_{1}}}{2\Omega_{1}}\,$, (22a)
$\dot{Y}_{1}$ $=-\left[\frac{\tilde{\Delta}^{2}_{01}}{2\Omega_{1}}+\frac{3\alpha}{4\Omega_{1}}(\frac{A_{1}^{2}}{2}+A_{2}^{2})\right]X_{1}-\frac{\Gamma}{2}Y_{1}+\frac{F_{1}\cos{\theta_{1}}}{2\Omega_{1}}\,$, (22b)
$\dot{X}_{2}$ $=-\frac{\Gamma}{2}X_{2}+\left[\frac{\tilde{\Delta}^{2}_{02}}{2\Omega_{2}}+\frac{3\alpha}{4\Omega_{2}}(\frac{A_{2}^{2}}{2}+A_{1}^{2})\right]Y_{2}+\frac{F_{2}\sin{\theta_{2}}}{2\Omega_{2}}\,$, (22c)
$\dot{Y}_{2}$ $=-\left[\frac{\tilde{\Delta}^{2}_{02}}{2\Omega_{2}}+\frac{3\alpha}{4\Omega_{2}}(\frac{A_{2}^{2}}{2}+A_{1}^{2})\right]X_{2}-\frac{\Gamma}{2}Y_{2}+\frac{F_{2}\cos{\theta_{2}}}{2\Omega_{2}}\,$, (22d)
where $A_{i}=\sqrt{X_{i}^{2}+Y_{i}^{2}}$ is the response amplitude at frequency $\Omega_{i}$, in analogy with Eq. (9). The nonlinearity introduces self-Kerr terms ($\propto A_{i}^{2}$), where the resonance frequency shifts based on the response at the tone itself, and cross-Kerr terms ($\propto A_{j}^{2}$), where the response at one tone shifts the effective resonance frequency of the other. These cross-terms capture the non-degenerate four-wave mixing interactions that were averaged out in the single-tone slow-beating approximation. By balancing harmonics only at $\Omega_{1}$ and $\Omega_{2}$ to obtain Eqs. (22), we neglect higher harmonics, e.g., ($3\Omega_{i}$, $5\Omega_{i}$) and mixed tones ($2\Omega_{1}-\Omega_{2}$, $2\Omega_{2}-\Omega_{1}$, …).
Solving for the stationary states of Eqs. (22) presents a significant challenge: the system comprises of four coupled cubic polynomials with a theoretical upper limit of 81 solutions [55]. Standard local root-finding algorithms risk missing coexisting stable states in such a complex landscape. To overcome this, we use the HarmonicBalance.jl package [53], which relies on homotopy continuation [56]. This approach allows us to globally map the solution space, rigorously tracking the number and stability of all stationary states as we sweep the control parameters $h$ and $\Delta_{21}$.
We first study the effect of the relative drive strength $h$ for a fixed detuning $\Delta_{21}$. In Fig. 7, we plot the stationary response amplitudes $A_{1}$ and $A_{2}$ (at frequencies $\Omega_{1}$ and $\Omega_{2}$, respectively) as a function of $h$, assuming the system is initialized in the low-amplitude attractor. For $h=0$, the system reduces to the single-tone case and the amplitude at $\Omega_{2}$ is 0. For positive detuning $\Delta_{21}>0$, see Fig. 7(a)], the response at $\Omega_{2}$ increases steadily with $h$, while the response at $\Omega_{1}$ remains initially confined to the low-amplitude state (blue branch in the $A_{1}$ panel). However, as the energy injected by the second tone increases with $h>0.25$, the amplitude of this lower branch at $\Omega_{1}$ rises rapidly until the branch destabilizes via a saddle-node bifurcation at $h=0.65$. Beyond this critical threshold, the system transitions to the high-amplitude branch (pink branch in the $A_{1}$ panel) at $\Omega_{1}$, which remains nearly constant, while the response at $\Omega_{2}$ continues to increase with $h$.
The stationary states presented in Fig. 7(a) reveal the mechanism preventing the secondary tone from becoming the dominant spectral component, despite its proximity to the bare resonance. Initially, the system exhibits a linear response regime where $A_{2}$ grows proportionally to the drive strength $h$, corresponding to the stable dynamics of Region 1. However, as the secondary amplitude becomes significant, the cross-Kerr interaction exerts a nonlinear frequency shift that pushes the tone’s effective resonance frequency away from $\Omega_{2}$ (towards lower frequencies for $\alpha<0$). This interaction creates a scenario reminiscent of an avoided crossing between the responses at $\Omega_{1}$ and $\Omega_{2}$: rather than $A_{2}$ capturing the input energy and increasing, beyond $h=0.25$ it plateaus and the renormalization of the resonance frequency redistributes power to pump $A_{1}$, the response at $\Omega_{1}$ (Region 2). Consequently, the nonlinearity “protects” the dominance of the primary response $A_{1}$, as the growing secondary tone pushes the effective resonance further towards $\Omega_{1}$. Ultimately, this cross-Kerr induced pumping drives $A_{1}$ to a critical threshold where the low-amplitude branch destabilizes via a saddle-node bifurcation, forcing the transition to the high-amplitude attractor observed in Region 3.
For negative detuning $\Delta_{21}<0$, see Fig. 7(b)], the stationary states display a distinct competition compared with the positive detuning case. The low-amplitude branch at $A_{1}$ and $A_{2}$ (solid blue line) remains stable over a wide range of $h$, with the primary response $A_{1}$ dominating. In contrast, the high-amplitude branch at $A_{1}$ and $A_{2}$ (solid pink line) exhibits a crossover where the secondary response $A_{2}$ grows rapidly, surpassing $A_{1}$ near $h\approx 0.5$, shortly before the branch destabilizes. As $h\rightarrow 1$, a parameter window emerges where no stable states exist at the ansatz frequencies.
This behavior correlates directly with the transition thresholds discussed in Fig. 5. The low-amplitude branch at $A_{1}$ and $A_{2}$ persists because the secondary tone is far-detuned from the bare resonance; it grows slowly and exerts negligible cross-Kerr influence on the primary mode. Conversely, for the high-amplitude state, the softening nonlinearity pulls the effective resonance frequency down, closer to the negatively detuned secondary tone. This reduces the effective detuning, enhancing the interaction through cross-Kerr terms and causing the branch to exhibit a large $A_{2}$. The crossover occurs in the vicinity of the blue/black transition threshold lines in Fig. 5; subsequently, the high-amplitude branch destabilizes, leaving the low-amplitude branch as the only available stable state. The absence of any stable states beyond $h=0.775$ indicates that the slow-flow variables $X_{i},Y_{i}$ never settle to a fixed point but instead oscillate. This corresponds to a regime of limit cycle dynamics [57], manifesting physically as persistent, self-sustained amplitude modulations.
To map the global stability landscape, we repeat the process above as a function of $h$ and $\Delta_{21}$, counting the number of stable branches, see Fig. 8. The black horizontal lines ($\Delta_{21}\sim\Gamma$) represent the limit of spectral resolution of the tones; the two-tone HBM is robust for well-resolved tones (i.e, $|\Delta_{21}|>\Gamma$). For positive detunings ($\Delta_{21}>0$), the numerically determined boundary where the system transitions from bistability (two stable branches, pink region) to monostability (one stable branch, green region) aligns remarkably well with the transition threshold derived from the high-amplitude renormalized filtered effective drive model (solid pink line). This confirms that, in the positively detuned regime ($\Delta_{21}>0$), the transition is fundamentally governed by a strong enough drive overcoming the resonance frequency shift induced by the target state itself.
In contrast, the negative detuning region ($\Delta_{21}<0$) reveals the breakdown of the stationary two-tone ansatz, consistent with the complex competition observed in Fig. 7(b). Here, the bifurcation boundaries diverge from the effective drive transition thresholds, indicating that the cross-Kerr interaction with the negatively detuned secondary tone induces parametric instabilities not captured by the simple resonance frequency shift model. We also observe that a distinct region emerges at high $h$ where the number of stable stationary branches drops to zero (purple region). This absence of stable fixed points provides strong evidence for the formation of limit cycles, confirming that the dynamical phase transitions in the negatively detuned regime involve time-dependent non-stationary states rather than simple inter-attractor switching.
Figure 8: The number of stable solutions plotted against varying $\Delta_{21}$ and $h$ for fixed $\Omega_{0}=2\pi$, $\Gamma=6.2\times 10^{-4}$, $F_{1}=4.81\times 10^{-4}$, and $\Delta_{10}=-9.02\times 10^{-4}$. The black lines at $\Delta_{21}=\pm 0.85\Gamma$ reflect the cuts shown in Fig. 7. 1 - 3 represent the fixed parameters used for the trajectories in Figs. 3 and 6. The colored transition thresholds shown in Fig. 5 have been superposed for comparison. A region with zero stable fixed points (purple) emerges, indicative of limit cycle dynamics.
The multifrequency ansatz [cf. Eq. (21)] and the resulting equations of motion [cf. Eq. (22)] successfully capture the cross-Kerr interaction (non-degenerate four-wave mixing) between the system’s responses to the two drives, providing physical insight beyond the linear and slow-beating approximations. The accuracy of this approach is confirmed in Fig. 6, where the HBM predictions at parameters 1- 3 align well with the distinct $\Omega_{1}$ and $\Omega_{2}$ components extracted from the Fourier spectra of the time-domain trajectories. However, the method relies on the assumption of slowly varying envelopes, which is strictly valid only when the tones are spectrally resolved ($|\Delta_{21}|\gtrsim\Gamma$). As the system approaches the slow-modulation limit ($\Delta_{21}\to 0$), this timescale separation breaks down. This leads to deviations between the multiple sharp bifurcation thresholds predicted by the HBM and the transitions observed in the phase diagram (cf. Fig. 5), which are further broadened by the coarse-graining inherent to the time-averaged order parameter $\bar{A}$. Finally, remaining discrepancies—particularly in the negative detuning regime—suggest that higher-order intermodulation products (e.g., $2\Omega_{1}-\Omega_{2}$) neglected by the two-tone truncation may play a non-negligible role in the dynamics.
VIII Conclusion/Outlook
In this work, we have developed a comprehensive framework to describe dynamical phase transitions in a Duffing resonator under bichromatic driving. By employing two complementary analytical approaches—an effective nonlinear response theory and an extended multifrequency ansatz—we have moved beyond phenomenological observation to uncover the distinct mechanisms governing the transitions in this system. Our analysis reveals that even a weak secondary tone can profoundly reshape the dynamics, not merely as a perturbation, but by fundamentally altering the stability landscape.
Our effective response theory provides a predictive tool for the slow-beating regime, where the secondary tone acts as a modulation that pushes the system past bifurcation points. Crucially, we identify that the onset of inter-attractor switching is not governed solely by the initial state, but by the resonance properties of the target state. Complementing this, our extended two-tone ansatz, valid in the resolved-sideband regime ($|\Delta_{21}|\gtrsim\Gamma$), reveals the microscopic mechanism behind these shifts. It reveals how cross-Kerr interactions (non-degenerate four-wave mixing) can either suppress switching via an avoided-crossing-like effect or induce instabilities leading to limit cycles, particularly in the negatively-detuned ($\Delta_{21}<0$) regime.
In contrast to earlier studies that surveyed specific parameters for sensing applications [29] or focused on cascades to chaos [30], we provided a rigorous mapping of the instability thresholds across a wide parameter space. This allowed us to explain the marked asymmetry between negative ($\Delta_{21}<0$) and positive ($\Delta_{21}>0$) detuning observed in experiments. Our results thereby offer a qualitative yet predictive framework to categorize dynamical phase transitions across nanomechanical [32], optical [37], and superconducting resonator [58, 38] platforms.
Looking ahead, this framework lays the foundation for systematic extensions. The multifrequency approach can be expanded to include higher-order intermodulation products, analytically clarifying the route to chaos in multi-tone driven systems. Furthermore, our stability maps provide a basis for investigating how stochastic or quantum fluctuations mediate transitions near these newly identified thresholds [59]. Finally, in the context of strongly nonlinear Kerr resonators, these insights into controlled state-switching offer new protocols for quantum control and qubit operation [60, 39], establishing multi-tone driving as a versatile resource for engineering dynamical phases.
@@ Acknowledgements
We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG) through project numbers 449653034, 521530974, and 545605411, as well as via SFB 1432 (project number 425217212). We also acknowledge support from the Swiss National Science Foundation (SNSF) through the Sinergia Grant No. CRSII5_206008/1. J.d.P. acknowledges funding from the Ramón y Cajal program (RYC2023-043827-I), funded by MICIU/AEI (10.13039/501100011033) and FSE+, and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M).
@@ Appendix B Exploration of more dynamical cases
Figure 10: The average amplitude phase diagram with threshold lines [cf. Fig. 5(c)], for an initial point starting from the higher branch, for $F_{1}\sim F_{lb}$. The parameters used are $\Omega_{0}=2\pi$, $\Delta_{10}=-9.02\times 10^{-4}$, $\Gamma=6.2\times 10^{-4}$, and $F_{1}=6.01\times 10^{-4}$.
@@ Appendix C Experimental platform
$\ddot{x}+\Omega_{n}^{2}x+\Gamma_{n}\dot{x}=F_{\mathrm{dr}}(t),$ (29)
where $\Omega_{n}$ and $\Gamma_{n}$ are the resonance frequency and the damping rate of the $n$-th mode, respectively, and $F_{\mathrm{dr}}(t)=\text{K}U(t)$ is the applied drive in units of $\mathrm{V}$\,$\mathrm{s}$^{-2}$. K is a conversion factor expressed in units of $\mathrm{s}$^{-2}$. When the strength of the external driving force increases, the resonator behavior starts to deviate from the standard damped harmonic oscillator. As we enter such large amplitude oscillation limit, to describe the system response to an external driving force we need to use the Duffing equation:
$\ddot{x}+\Omega_{n}^{2}x+\Gamma_{n}\dot{x}+\alpha_{n}x^{3}=\text{Re}\left[F_{\mathrm{dr}}e^{i(\Omega_{\mathrm{dr}}t+\theta)}\right],$ (30)
@@
In this work, we only focus on the lowest mechanical mode of the resonator. The mechanical parameters for this particular device have been characterized in a previous work [31] by measuring the amplitude ($A$) and phase ($\phi$) response to an external drive in both the weak and strong drive limit. Using a lock-in amplifier, we apply an external driving force $\Omega_{dr}$ to the system, varying its value across the resonance $\Omega_{0}$. With the same instrument, we simultaneously measure $A$ and $\phi$ in the frame rotating at $\Omega_{dr}$. The measured mechanical parameters are $\Omega_{0}/2\pi=1.11\,$\mathrm{M}\mathrm{H}\mathrm{z}$\$, $\Gamma/2\pi=108\,$\mathrm{H}\mathrm{z}$\$, $\alpha=-1.89\,$\mathrm{V}$^{-2}\,\mathrm{s}^{-2}$ and $\text{K}\approx 1\times 10^{7}s^{-2}$.
To generate the phase diagram in Fig. 5(a), we measure the resonator amplitude $A$ as a function of time when driven with two different tones $F_{1}e^{i(\Omega_{1}t)}$ and $F_{2}e^{i(\Omega_{2}t))}$. We use the lock-in amplifier to apply the two driving tones and readout the amplitude in the frame rotating at $\Omega_{1}$. After fixing $F_{1}/\text{K}=140\mathrm{mV}$ and $\Delta_{10}/2\pi\approx-158\,{Hz}$, we systematically probe the response of the resonator while varying $\Delta_{21}=\Omega_{2}-\Omega_{1}$ and $h$. For each combination, we extract the average value $\bar{A}$ for an interval of $5\,$\mathrm{s}$\$ making sure that for each combination the system is initialized in the lower stable solution of the Duffing curve before turning on the second drive. An example of measured $A(t)$ in the two limits of $\Delta_{21}<\Gamma$ and $\Delta_{21}\approx\Gamma$ are shown in Fig. 13.